At a distance of 0.10 m the potential of the point like charge is 75 V. At what is the potential equivalent to 25 V?
Answer #1

Formula for potential is

75=kQ/0.10 ...... equation (1)

We can write, potential as

25 = kQ/R ....... equation (2)

on dividing equation (1) by equation (2)  

=> 3=R/0.10

R=3*0.10

R=0.3 m (Answer)

I hope help you !!